OpenStudy (anonymous):

Solve the system using elimination. 5x + 8y = -29 7x - 2y = -67

OpenStudy (anonymous):

HELP!

OpenStudy (anonymous):

Multiply both sides of the second equation by 4 and then add the equations together. +8y -8y will subtract out leaving one equation in one unknown, x.

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

solving by elimination you are trying to cancel out one term with the single remaing term x - y = 0 x - y = 0 solve for y x - y = 0 -- multiply to cancel out 'x or y' + x - y = 0 -- I chose 'x' ------------ y = x use the variable you solved for substitute it back in to x - y = 0 <----- if you cancelled out the with this (leaving one variable alone) x - y = 0 <--- substitute it with this formula 'y' value x - y = 0 <---substitute it with this formula 'y' value x - y = 0 <-------if you cancelled out with this you should get your then get your x result y result (what you used to substitute with)